OpenStudy (anonymous):

When solving negative one over three.(x − 15) = −4, what is the correct sequence of operations?

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

\[-\frac{1}{3}\left(x-15\right)=-4\]?

OpenStudy (texaschic101):

well....you can either distribute the -1/3 through the parenthesis...or...you can multiply both sides by 3

OpenStudy (anonymous):

So, first i would multiply each side by -3 and add 15 to each side? @texaschic101